# GRADA: Graph-based Reranker against Adversarial Documents Attack

### v1

@@ Abstract
Retrieval Augmented Generation (RAG) frameworks improve the accuracy of large language models (LLMs) by integrating external knowledge from retrieved documents, thereby overcoming the limitations of models’ static intrinsic knowledge. However, these systems are susceptible to adversarial attacks that manipulate the retrieval process by introducing documents that are adversarial yet semantically similar to the query. Notably, while these adversarial documents resemble the query, they exhibit weak similarity to benign documents in the retrieval set. Thus, we propose a simple yet effective Graph-based Reranking against Adversarial Document Attacks (GRADA) framework aiming at preserving retrieval quality while significantly reducing the success of adversaries. Our study evaluates the effectiveness of our approach through experiments conducted on five LLMs: GPT-3.5-Turbo, GPT-4o, Llama3.1-8b-Instruct, Llama3.1-70b-Instruct, and Qwen2.5-7b-Instruct. We use three datasets to assess performance, with results from the Natural Questions dataset demonstrating up to an 80% reduction in attack success rates while maintaining minimal loss in accuracy.
GRADA: Graph-based Reranker against Adversarial Documents Attack
Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3, Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu5 1University of Melbourne, 2University of Edinburgh, 3University College London 4University of Manchester, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@manchester.ac.uk, qiongkai.xu@mq.edu.au
@@ 1 Introduction
While RAG enhances the flexibility of LLMs, it also introduces new vulnerabilities. Adversaries can exploit retrieval mechanisms by injecting manipulated documents into the corpus Zhong et al. (2023); Clop and Teglia (2024); Greshake et al. (2023); Pasquini et al. (2024), subtly altering rankings to mislead LLM outputs. As shown in Figure 1, these adversarial documents mimic query-relevant patterns, making them difficult to detect while degrading the reliability of retrieval-based LLM systems. In practice, LLMs are used in the search engine to produce direct answers to user’s query, called Answer Engine Optimization (AEO)  Yalçın and Köse (2024). Adversarial documents can be created using the previous methods to achieve higher rankings in the search results. Also, guide the search engine to produce malicious context that can potentially harm the user Hammond (2024); Venkit et al. (2024).
@@
Figure 2: An overview of GRADA. A vanilla RAG pipeline concatenates all retrieved documents along with the question as the input to the LLM. However, the accuracy of this pipeline can be easily harmed by malicious passages. In contrast, GRADA uses a graph-based approach to isolate and filter out malicious passages before passing the retrieved documents as the LLM input.
@@
Existing noise filtering methods, such as Hybrid List Aware Transformer Reranking (HLATR, Zhang et al., 2022) and BAAI General Embeddings (BGE-reranker, Xiao et al., 2023), focus on improving document relevance by filtering out generic noise or low-quality content. However, these methods are ineffective against adversarial attacks that exploit query-document similarity patterns to evade detection. In addition, a recent study has reviewed current graph-based reranking methods. It shows a potential path of using graphs in future information retrieval Zaoad et al. (2025) tasks but the effects on adversarial documents remains unknown. On the other hand, specialized adversarial defenses, such as keyword filtering and decoding aggregation Xiang et al. (2024), can successfully remove adversarial content but at the cost of discarding valuable benign documents, ultimately weakening retrieval performance. This trade-off highlights the need for a more nuanced defense mechanism that can distinguish between adversarial and benign documents without compromising retrieval quality.
@@
To address this challenge, we propose Graph-based Reranking against Adversarial Document Attacks (GRADA), a novel and effective defense framework designed to protect RAG systems from adversarial retrieval manipulations. Our key insight is that adversarial documents, while optimized for high query similarity, exhibit weaker semantic coherence with genuinely relevant documents in the retrieval set. Leveraging this property, we construct a graph where each retrieved document is represented as a node, and edges capture document-document similarity relationships. By propagating ranking scores through this graph structure, our approach prioritizes clusters of semantically consistent documents while suppressing adversarially crafted outliers. As illustrated in Figure 2, our method significantly enhances the robustness of RAG-based LLMs, mitigating adversarial influences while preserving the integrity of benign retrieval results.
@@
We conducted comprehensive experiments on Natural Questions (NQ), MS-MARCO, and HotpotQA across five different models. Our method has shown at least a 30% decrease in reducing the Attack Success Rate (ASR), with improvements of up to 80% across various adversarial attack strategies.
@@
  • We introduce GRADA, a weighted similarity graph among retrieved documents iteratively propagates scores to mitigate the impacts of adversarial passages.
  • We introduce a novel scoring function that simultaneously captures both query-document and document-document correlations, thereby improving robustness against adversarial attempts to mimic the query.
  • Conducted comprehensive experiments across three different datasets against four chosen attacks. Showing GRADA’s advantages over the current defense baselines.
@@ 2 Related Work
The IR community has long examined how malicious actors distort ranking functions. Gyongyi and Garcia-Molina (2005) propose a taxonomy of web-spam strategies, categorizing attacks into content-based, link-based, and behavior-based manipulations. Content-based spam involves manipulating the textual content of web pages to boost rankings, such as stuffing keywords or inserting irrelevant but popular terms. Link-based spam, including techniques like link farms, exploits the graph structure of the web to artificially inflate a page’s authority. Ntoulas et al. (2006) focus on detecting such content-based spam using statistical features extracted from term distributions and document structure. Their work highlights how spam pages often exhibit unnatural language models and disproportionately high keyword densities. Castillo and Davison (2011) provide a comprehensive survey of adversarial techniques in web search, including cloaking (serving different content to users and crawlers), redirection, and deceptive anchor texts. These traditional adversarial IR strategies, although originally developed to mislead search engines, reveal core vulnerabilities in retrieval systems—vulnerabilities that persist, albeit in more complex forms, in modern neural IR and RAG systems.
@@
When RAG systems came out, Corpus poisoning attacks Zhong et al. (2023) show a possible new attack surface on LLMs. However, this method does not directly affect the accuracy of the LLM; instead, it focuses on the retriever. Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al. (2023); Pasquini et al. (2024). However, compared to the prior work, these methods are unstable in ensuring the retriever retrieves the adversarial passage every time. While these attacks are situated in modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
@@
More recently, PoisonedRAG Zou et al. (2024) was proposed as a more stable attack. It uses two passages concatenated together, with one of them appended to guarantee the retrieval of the adversarial passage and one to achieve a given adversarial goal on the generator. The goal is to let the LLM output the answer the attacker wants. PoisonedRAG inspired a lot of the new attacks. Phantom Chaudhari et al. (2024), which introduces a trigger to the question and achieves the adversarial goal only when the trigger is shown in the query. Another Prompt Injection Attack (PIA, Clop and Teglia, 2024) makes use of the passage that guarantees the retrieval in PoisonedRAG and focuses on the adversarial goal beyond misinformation.
A recent study proposed a defense mechanism that generates responses independently and produces an output based on the majority vote Xiang et al. (2024). However, this method initiates its defense at the generator stage, which can impact the accuracy of the system, especially when multiple documents are required. GRADA addresses this issue by focusing on the stage before generation, specifically the reranking process.
@@
Figure 3: BM25 similarity matrix among retrieved documents, where D0-D4 are poisoned, and D5-D10 are clean.
@@ 3 GRADA
A defining characteristic of recent poisoning attacks on RAG Zou et al. (2024) is their focus on ensuring semantic similarity to the query while introducing anomalous similarity patterns among poisoned documents. Adversarial documents are very similar to the query but less to the legitimate documents, forming isolated patterns within the retrieval set, as illustratedFigures 2 and 3. Graph structures naturally capture these complex inter-document relationships by representing documents as nodes and similarities as edges. Leveraging this intuition, we propose a graph-based reranking method that utilizes document-document similarity to enhance retrieval robustness. In Section 3.1, we detail the graph construction process, followed by a description of our reranking system in Section 3.2.
@@ 3.1 Graph Construction
We construct a weighted, undirected graph $G=(V,E)$, where each node $v_{i}\in V$ corresponds to a document $doc_{i}$, and each edge $e_{ij}\in E$ is an undirected edge connecting node $v_{i}$ and $v_{j}$. Each edge is assigned a weight $w_{ij}\in\mathbb{R}^{+}$, which quantifies the similarity between the corresponding documents, i.e., $\text{sim}(v_{i},v_{j})$. The graph is undirected because document relationships are not inherently directional; rather, the connectivity structure defines their associations. The edge weight $w_{ij}$ can be computed using two different approaches:
@@
    $w_{ij}=f\big{(}\text{sim}(v_{i},v_{j}),\text{sim}(v_{i},q),\text{sim}(v_{j},q)\big{)}$
@@
$\text{sim}_{\text{sum}}=\text{sim}(v_{i},q)+\text{sim}(v_{j},q)$
@@
$w_{ij}=\max\left(\text{sim}(v_{i},v_{j})-\alpha\cdot\text{sim}_{\text{sum}},0\right)$
Here, $\alpha$ is a penalty coefficient that controls the influence of query similarity. If $\text{sim}(v_{i},v_{j})<\alpha\cdot\left[\text{sim}\left(v_{i},q\right)+\text{sim}\left(v_{i},q\right)\right]$, the edge weight is set to zero, effectively removing the connection between $v_{i}$ and $v_{j}$.
@@
    $w_{ij}=\frac{1}{2}\left(\text{BM25}\left(v_{i},v_{j}\right)+\text{BM25}\left(v_{j},v_{i}\right)\right)$
@@
    $w_{ij}=\text{sim}(v_{i},v_{j})=\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|}$
@@ 3.2 Reranking
$s_{i}^{(t)}$ $=(1-d)s^{*}_{j}+d\sum_{v_{j}\in\mathcal{N}(i)}\frac{w_{ij}}{\sum_{v_{k}\in\mathcal{N}(j)}w_{jk}}s_{j}^{(t-1)}$ (1)
where $N(i)$ represents the set of neighbor nodes connected by $v_{i}$ and $d$ is the damping factor, typically set to 0.85. The initial score vector $s^{*}$ is set by uniform initialization $s^{*}=\left[\frac{1}{|V|},\frac{1}{|V|},...,\frac{1}{|V|}\right]$.
For experiments comparing different initialization methods, please refer to Section C.2.
The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval and $M\geq n$. We retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
By ensuring that poisoned documents do not form the majority in the retrieved set (with $M\geq 2n$), we prevent adversarial documents, which may exploit the query for high relevance scores, from dominating. For example, if the original set of $n$ documents contains all poisoned ones (e.g., $n=5$), adding $\geq n$ benign documents ensures the majority is non-poisoned. This strategy guarantees that non-poisoned documents remain a significant portion of the final selection, enhancing the system’s robustness against adversarial manipulation.
After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded. These top $n$ documents are then provided as the context of the model.
@@ 4.1 Experimental Setup
We conduct experiments on three widely used english datasets: Natural Question Kwiatkowski et al. (2019), MS-MARCO Nguyen et al. (2016) and HotpotQA Yang et al. (2018). The victim models chosen for this study are GPT-3.5-Turbo (version 0125) Brown et al. (2020), GPT-4o (version 2024-08-06) OpenAI et al. (2024), Qwen2.5 Qwen et al. (2025) and LLaMA-3 Grattafiori et al. (2024). The prompts used to generate answers are detailed in Appendix A. Contriever Izacard et al. (2021), is a dense retriever model used to find relevant documents by calculating similarity scores between the query and the knowledge base. It was selected for this study due to its efficiency and ability to handle large datasets. In this work, we investigate four distinct attack strategies on RAG. Two of them are Black-box attacks that have no knowledge about the retriever: PoisonedRAG Zou et al. (2024) and PIA Greshake et al. (2023); Pasquini et al. (2024); Perez and Ribeiro (2022). The remaining two are white-box attacks, in which the attacker has access to the victim’s retriever: PoisonedRAG(Hotflip) Zou et al. (2024) and Phantom Chaudhari et al. (2024)
@@
However, in contrast to Zou et al. (2024), where five poisoned texts are generated and injected into the knowledge base, To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database. The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack. As shown in Figure 3, a similarity cluster of poisoned documents appears in the top-left corner. By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries. This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
Defense setup.
We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function. [^1] We examine other similarity functions in Section 4.3 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024). We compare GRADA against two reranking models and one defense method: HLATR Zhang et al. (2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al. (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al. (2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section 3. We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation). The top five ranked documents are subsequently provided as the context for the model to answer the query. This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question. For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
Evaluation metrics.
To assess the question-answering accuracy of the models, we adopt EM score. EM requires that the predicted answer of the model matches the ground truth answer exactly. This strict criterion ensures that the response of the model is precise and follows the need for exact wording specified in the query, as outlined in Appendix A.
@@
Table 1: ASR and EM (%) for various defense methods on the black-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct. The results of other models can be found in Tables 10, 9, 11, 13 and 12. We highlight the top-2 lowest ASR results in blue cells.
@@
4.2 Results and Discussions
Attacking without defense.
As shown in Table 1, including a single poisoned document in the retrieval process results in a high ASR score. For instance, PoisonedRAG achieves an ASR of 50% across three datasets on both GPT-3.5-Turbo and Llama3.1-8b-Instruct. PIA achieves at least 69% ASR on Llama3.1-8b-Instruct and up to 100% ASR in GPT-3.5-Turbo. These findings emphasize that even minimal adversarial input can achieve very high ASR and degrade the model’s accuracy.
Effectiveness of GRADA.
The impact of GRADA on mitigating adversarial attacks is demonstrated in Tables 1 and 2. As shown in Table 1, on the NQ and MS-MARCO datasets using GPT-3.5-Turbo, the ASR for PIA decreases from 98.0% and 88.0% to 2.0% and 3.0% by using D2DSIM-EBD. With D2DSIM-EBD, GRADA is also effective against PoisonedRAG, effectively reducing the ASRs from 55.7% and 46.5% to 26.1% and 29.0%. However, the reduction of ASR against PoisonedRAG is more modest than against the other attacks. On this attack, D2DSIM-BM25 and HRSIM led to significant improvements compared to D2DSIM-EBD, where D2DSIM-BM25 achieved an extra 13% decrease in ASR to 13.5% and 16.5%. Beyond that, HRSIM which introduces penalties for excessive similarity to the query, finalizes the ASR to 3% and 8.5%.
In Table 1, HLATR and BGE-reranker exhibit limited ability to filter poisoned documents, with ASR remaining largely unchanged compared to scenarios without any defense mechanisms. Although BGE-reranker occasionally outperforms HLATR, its overall performance remains inferior to GRADA in handling adversarial cases. This discrepancy underscores a critical limitation in contemporary reranking systems, which are primarily optimized for question relevance but insufficiently equipped to address adversarial attacks with high question relevance.
Keyword Aggregation is able to reduce ASR significantly, especially for attacks like PIA and Phantom. Keyword Aggregation works by extracting keywords from the answers of each passage to generate the final response, effectively neutralizing attack payloads designed to manipulate or deny answers, such as producing advertisements. However, while it reduces ASR effectively, its EM scores are lower than those of GRADA. For example, on Llama3.1-8b-Instruct in Table 1, GRADA’s EM scores dominate Keyword Aggregation with at most 21% difference as some critical information may be lost during keyword extraction. This shows the ability of GRADA to perform well on normal answers even after mitigating adversarial contents.
Similar results to those presented in Table 1 can also be observed in Table 2. Notably, GRADA combined with HRSIM consistently outperforms all other approaches, demonstrating that HRSIM is a strong similarity scoring function compared to the alternatives used in GRADA.

### v2

# GRADA: Graph-based Reranking against Adversarial Documents Attack

@@ Abstract
Retrieval Augmented Generation (RAG) frameworks can improve the factual accuracy of large language models (LLMs) by integrating external knowledge from retrieved documents, which is useful for overcoming the limitations of models’ static intrinsic knowledge. However, these systems are susceptible to adversarial attacks that manipulate the retrieval process by introducing documents that are adversarial yet semantically similar to the query. Notably, while these adversarial documents resemble the query, they exhibit weak similarity to benign documents in the retrieval set. Thus, we propose a simple yet effective Graph-based Reranking against Adversarial Document Attacks (GRADA) framework aimed at preserving retrieval quality while significantly reducing the success of adversaries. Our study evaluates the effectiveness of our approach through experiments conducted on six LLMs: GPT-3.5-Turbo, GPT-4o, Llama3.1-8b-Instruct, Llama3.1-70b-Instruct, Qwen2.5-7b-Instruct, and Qwen2.5-14b-Instruct. We use three datasets to assess performance, with results from the Natural Questions dataset showing up to an 80% reduction in attack success rates while maintaining minimal loss in accuracy.
GRADA: Graph-based Reranking against Adversarial Documents Attack
Jingjie Zheng1, Aryo Pradipta Gema2, Giwon Hong2, Xuanli He3thanks: Corresponding Author., Pasquale Minervini2, Youcheng Sun4, Qiongkai Xu1,511footnotemark: 1 1University of Melbourne, 2University of Edinburgh, 3University College London 4Mohamed bin Zayed University of Artificial Intelligence, 5Macquarie University jingjzheng@student.unimelb.edu.au, zodiachy@gmail.com {aryo.gema, giwon.hong, p.minervini}@ed.ac.uk youcheng.sun@mbzuai.ac.ae, qiongkai.xu@mq.edu.au
@@ 1 Introduction
While RAG enhances the flexibility of LLMs, it also introduces new vulnerabilities. Adversaries can exploit retrieval mechanisms by injecting manipulated documents into the corpus Zhong et al. (2023); Clop and Teglia (2024); Greshake et al. (2023); Pasquini et al. (2024), subtly altering rankings to mislead LLM outputs. As shown in Figure˜1, these adversarial documents mimic query-relevant patterns, making them difficult to detect while degrading the reliability of retrieval-based LLM systems. In real-world applications, LLMs are increasingly used in search engines to provide direct answers to user queries, a process known as Answer Engine Optimization (AEO) Yalçın and Köse (2024). By leveraging retrieval-time manipulation techniques, attackers can craft adversarial content that not only ranks higher in search results but also steers the generated answers toward harmful or misleading content Hammond (2024); Venkit et al. (2024).
@@
Figure 2: An overview of GRADA. A vanilla RAG pipeline concatenates all retrieved documents along with the question as input to the LLM. However, the accuracy of this pipeline can be easily harmed by malicious passages. In contrast, GRADA uses a graph-based approach to rerank and filter out malicious passages before passing the retrieved documents to LLMs for generation.
@@
Existing noise filtering methods, such as Hybrid List Aware Transformer Reranking (HLATR, Zhang et al., 2022) and BAAI General Embeddings (BGE-reranker, Xiao et al., 2023), focus on improving document relevance by filtering out generic noise or low-quality content. However, these methods are ineffective against adversarial attacks that exploit query-document similarity patterns to evade detection. In addition, a recent study has reviewed current graph-based reranking methods Zaoad et al. (2025). It shows a potential path to use graphs in future information retrieval tasks, but the effects on adversarial documents remain unknown. On the other hand, specialized adversarial defenses, such as keyword filtering and decoding aggregation Xiang et al. (2024), can successfully remove adversarial content but at the cost of discarding valuable benign documents, ultimately weakening retrieval performance. This trade-off highlights the need for a more nuanced defense mechanism that can distinguish between adversarial and benign documents without compromising retrieval quality.
@@
To address this challenge, we propose Graph-based Reranking against Adversarial Document Attacks (GRADA), an effective defense framework designed to protect RAG systems from adversarial retrieval manipulations. Our key insight is that adversarial documents, while optimized for high query similarity, exhibit weaker semantic coherence with genuinely relevant documents in the retrieval set. Leveraging this property, we construct a graph where each retrieved document is represented as a node, and edges capture document-document similarity relationships. By propagating ranking scores through this graph structure, our approach prioritizes clusters of semantically consistent documents while suppressing adversarially crafted outliers. As illustrated in Figure˜2, our method significantly enhances the robustness of RAG-based LLMs, mitigating adversarial influences while preserving the integrity of benign retrieval results.
@@
We conducted comprehensive experiments on Natural Questions (NQ), MS-MARCO, and HotpotQA across six different models. Our method has shown at least a 30% decrease in reducing the Attack Success Rate (ASR), with improvements of up to 80% across various adversarial attack strategies.
@@
  • We introduce GRADA, which constructs a weighted similarity graph among retrieved documents and iteratively propagates scores to mitigate the impacts of adversarial passages.
  • We introduce a novel scoring function that simultaneously considers both query-document and document-document correlations, thereby improving robustness against adversarial attempts to mimic the query.
  • We conducted comprehensive experiments on three distinct datasets, evaluating our method against four representative attack types. The results consistently demonstrate that GRADA outperforms existing defense baselines.
@@ 2 Related Work
Adversarial manipulation in IR has a long history. Gyongyi and Garcia-Molina (2005) categorize web-spam strategies into content-based, link-based, and behavior-based attacks, while Ntoulas et al. (2006) use statistical features to detect spam content. Castillo and Davison (2011) survey a range of traditional attacks like cloaking and redirection, which expose fundamental weaknesses that persist in modern neural retrieval systems.
@@
When RAG systems came out, Corpus poisoning attacks Zhong et al. (2023) and third-party API attacks Zhao et al. (2024) show a new potential attack surface on LLMs. Later, prompt injection attacks were introduced to bypass the retriever and affect the generator successfully Greshake et al. (2023); Pasquini et al. (2024). However, compared to the prior work, these methods are unstable in the retrieved adversarial passages. While these attacks are based on modern LLM-based retrieval, adversarial manipulation of information-retrieval systems has a much longer history that is instructive for our setting.
@@
More recently, PoisonedRAG Zou et al. (2024) was proposed as a more stable attack. It uses two passages concatenated together, with one of them appended to guarantee the retrieval of the adversarial passage and one to achieve the adversarial goal on the generator, which is to steer the LLM generating the answers anticipated by the attacker. PoisonedRAG inspired many subsequent attacks. Phantom Chaudhari et al. (2024), which introduces a trigger to the question and achieves the adversarial goal only when the trigger is shown in the query. Another type of Prompt Injection Attack (PIA, Clop and Teglia, 2024) leverages the guaranteed retrieval mechanism in PoisonedRAG. Unlike typical misinformation attacks, this variant targets broader adversarial objectives beyond merely spreading false information.
A recent study proposed a defense mechanism that generates responses independently and produces an output based on the majority vote Xiang et al. (2024). While effective in some settings, this strategy defends only at the generator stage, which can compromise accuracy when multiple documents must be integrated. In contrast, several recent works intervene earlier in the RAG pipeline. TrustRAG Zhou et al. (2025) employs clustering and LLM-based conflict resolution to filter poisoned documents before they influence the generator, substantially reducing attack success rates while preserving accuracy. Complementary to this, traceback approaches such as RAGForensics Zhang et al. (2025) focus on identifying and removing the poisoned texts within the knowledge base itself, ensuring that subsequent retrieval yields only benign passages. Building on these insights, GRADA strengthens defenses at the reranking stage, preventing malicious content from reaching the generator without sacrificing the benefits of multi-document retrieval.
@@
Figure 3: The similarity matrix among retrieved documents using BM25, where D0-D4 are poisoned and D5-D10 are clean documents.
@@ 3 GRADA
A defining characteristic of recent poisoning attacks on RAG Zou et al. (2024) is their focus on ensuring semantic similarity to the query while introducing anomalous similarity patterns among poisoned documents. Adversarial documents closely resemble the query while diverging significantly from the legitimate documents, resulting in isolated patterns within the retrieval set, as illustrated in Figures˜2 and 3. Graph structures naturally capture these complex inter-document relationships by representing documents as nodes and similarities as edges. Leveraging this intuition, we propose a graph-based reranking method that utilizes document-document similarity to enhance retrieval robustness. In Section˜3.1, we detail the graph construction process, followed by a description of our reranking system in Section˜3.2.
@@ 3.1 Graph Construction
We construct a weighted, undirected graph $\mathcal{G}=(V,E)$, where each node $v_{i}\in V$ corresponds to a document, and each edge $e_{ij}\in E$ is an undirected edge connecting node $v_{i}$ and $v_{j}$. Each edge is assigned a weight $w_{ij}\in[0,1]$, which quantifies the similarity between the corresponding documents, i.e., $\text{sim}(v_{i},v_{j})$. The graph is undirected because document relationships are not inherently directional; rather, the connectivity structure defines their associations.
The edge weight $w_{ij}$ can be computed using different approaches as follows:
@@
    $w_{ij}=f\big{(}\text{sim}(v_{i},v_{j}),\text{sim}(v_{i},q),\text{sim}(v_{j},q)\big{)}.$
@@
$\text{sim}_{\text{sum}}=\text{sim}(v_{i},q)+\text{sim}(v_{j},q).$
@@
$w_{ij}=\max\{\text{sim}(v_{i},v_{j})-\alpha\cdot\text{sim}_{\text{sum}},0\}.$
Here, $\alpha$ is a penalty coefficient that controls the influence of query similarity. If $\text{sim}(v_{i},v_{j})<\alpha\cdot\left[\text{sim}\left(v_{i},q\right)+\text{sim}\left(v_{i},q\right)\right]$, the edge weight is set to zero, which indicates removing the collusion connection between $v_{i}$ and $v_{j}$.
@@
    $w_{ij}=\frac{1}{2}\left(\text{BM25}\left(v_{i},v_{j}\right)+\text{BM25}\left(v_{j},v_{i}\right)\right).$
@@
    $w_{ij}=\text{sim}(v_{i},v_{j})=\frac{\mathbf{x}_{i}\cdot\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|}.$
@@ 3.2 Reranking
$s_{i}^{(t)}$ $=(1-d)s^{*}_{j}+d\sum_{v_{j}\in\mathcal{N}(i)}\frac{w_{ij}}{\sum_{v_{k}\in\mathcal{N}(j)}w_{jk}}s_{j}^{(t-1)},$ (1)
where $\mathcal{N}(i)$ represents the set of neighbor nodes connected by $v_{i}$ and $d$ is the damping factor, typically set to 0.85.111The experiments comparing different initialization methods are provided in Section C.3. The initial score vector $s^{*}$ is set by uniform initialization $s^{*}=\left[\frac{1}{|V|},\frac{1}{|V|},...,\frac{1}{|V|}\right]$.
The framework works as follows: The retriever identifies $M$ documents most similar to the query, with $n$ being the number of documents originally intended for retrieval (usually $M\geq n$). In experiments, we retrieve additional documents to maintain consistency in the number of documents in the non-defended scenario.
To prevent adversarial documents from dominating the retrieved set, we ensure that poisoned documents do not constitute the majority by retrieving at least twice the number of documents (i.e., $M\geq 2n$). For instance, if all $n$ original documents are poisoned (e.g., $n=5$), incorporating at least $n$ additional benign documents guarantees that the majority of the final selection is non-poisoned. This approach maintains a substantial presence of benign content in the retrieved set, thereby improving the system’s resilience to adversarial manipulation.
After the algorithm reaches a stationary score distribution, the top $n$ documents are retained, while the remaining documents are discarded. Then, these top $n$ documents are provided as the context of the model.
@@ 4.1 Experimental Setup
We conduct experiments on three widely used English datasets: Natural Question Kwiatkowski et al. (2019), MS-MARCO Nguyen et al. (2016) and HotpotQA Yang et al. (2018). The victim models chosen for this study are GPT-3.5-Turbo (version 0125) Brown et al. (2020), GPT-4o (version 2024-08-06) OpenAI et al. (2024), Qwen2.5 Qwen et al. (2025) and LLaMA-3 Grattafiori et al. (2024). The prompts used to generate answers are detailed in Appendix˜A. Contriever Izacard et al. (2021), is a dense retriever model used to find relevant documents by calculating similarity scores between the query and the knowledge base. It was selected for this study due to its efficiency and ability to handle large datasets. In this work, we investigate four distinct attack strategies on RAG. Two of them are Black-box attacks that have no knowledge about the retriever: PoisonedRAG Zou et al. (2024) and PIA Greshake et al. (2023); Pasquini et al. (2024); Perez and Ribeiro (2022). The remaining two are white-box attacks, in which the attacker has access to the victim’s retriever: PoisonedRAG(Hotflip) Zou et al. (2024) and Phantom Chaudhari et al. (2024)
@@
In contrast to Zou et al. (2024), where five poisoned texts are generated and injected into the knowledge base. To provide a more realistic assessment of the attack’s effectiveness, we modify the experiment to inject only a single poisoned document into the database. The original setup, which retrieved only poisoned documents, resulted in a 100% Attack Success Rate (ASR), making it impractical to evaluate the true impact of the attack. As shown in Figure˜3, a similarity cluster of poisoned documents appears in the top-left corner. By applying a clustering algorithm, we can identify and merge redundant information, effectively removing repetitive poisoned entries. This adjustment ensures that only one poisoned document is retrieved, allowing for a more meaningful evaluation of the attack’s success.
@@
Table 1: ASR and EM (%) for various defense methods on the black-box attack methods on GPT-3.5-Turbo and Llama3.1-8b-Instruct. The results of other models can be found in Tables˜11, 10, 12, 14 and 13. We highlight the top-2 lowest ASR results in blue cells.
Defense setup.
We explore three similarity score combinations for GRADA: Embedding-based Distance, BM25, and Hybrid Relevance Similarity with BM25 as the similarity function. [^2] We examine other similarity functions in Section C.2 Here, we utilize Contriever to encode both documents and queries, while for BM25, we adopt the implementation provided by  Lù (2024). We compare GRADA against two reranking models and one defense method: HLATR Zhang et al. (2022), which achieved first place in the MS-MARCO Passage Ranking Leaderboard, BGE-reranker Xiao et al. (2023), which achieves a high precision score in ranking tasks, and Keyword Aggregation Xiang et al. (2024), the only existing defense specifically designed for RAG-based adversarial attacks, as a baseline.
We evaluate the effectiveness of these defense methods by integrating them into our two-stage retrieval system described in Section˜3. We initially retrieve $M=10$ documents, which are then reranked using the aforementioned methods (except for Keyword Aggregation). The top five ranked documents are subsequently provided as the context for the model to answer the query. This ensures that, regardless of the defense configuration, the model always receives a fixed number of five context documents to respond to the question. For Keyword Aggregation, which does not perform reranking, the model directly generates the output based on the algorithm’s selection.
Evaluation metrics.
To assess the question-answering accuracy of the models, we adopt EM score. EM requires that the predicted answer of the model matches the ground truth answer exactly. This strict criterion ensures that the response of the model is precise and follows the need for exact wording specified in the query, as outlined in Appendix˜A.